# Data Science Principles for Interpretable and Explainable AI

### v1

@@ 3 Evaluation
Operationalization: Interpretability encompasses several technical definitions, like parsimony and simulatability, making it difficult to capture precisely with a single metric.
@@
Ambiguous Ground Truth: Unlike supervised learning, establish ground truth interpretability output (e.g., the “correct” feature attribution map) is often impossible.
Human-Computer Interaction: The success of an interpretability method depends on how it interfaces with users. Simplified experiments may not reflect real-world complexity, while real-world observations may be confounded by factors that don’t generalize across applications.
@@
Interpretability evaluation methods fall into two general categories: tailored benchmark datasets and user studies with controlled tasks. Within each class, there are variations that account for the challenges above in different ways. For dataset benchmarks, one approach is ablation (Hooker et al., 2019; Wang et al., 2022). Specifically, for methods that output feature importance attributions, this involves retraining models versions of the data where purported important regions are masked out. Deterioration in model performance suggests that the attributions accurately reflect regions that the original model relied upon. Note that retraining is necessary. The ablated images are not drawn from the same distribution as the training data, so deterioration in performance of the original model could be due to either removal of important features (the criteria of interest) or domain shift (a confounding factor). Another data-driven benchmarking approach is to design synthetic datasets. These datasets are generated using known mechanisms, allowing for later evaluation of explanations. For example, the funnybirds benchmark creates imaginary cartoon birds where “species” classes are derived from attributes, like wing, beak, and tail type, that are controlled during data generation (Hesse et al., 2023). Given a specific instance, an ideal explanation should focus on the parts of the bird that distinguish it from related classes. To make these benchmarks more realistic, variation in viewpoint and illumination can be introduced.
@@
In user studies, one approach is to have participants make pairwise preference judgments between explanations (Jeyakumar et al., 2020). However, subjective preference may not align with objective task performance. To this end, more specific experimental tasks can be designed, evaluating user decisions with and without support from interpretability outputs (Ma et al., 2024; Colin et al., 2022). For example, to assess simulatability, participants could be asked to guess a model’s output on an example before and after viewing an explanation. Alternatively, they could be asked to guess how predictions might change under perturbations to the input – effective explanations should help in making accurate guesses. Finally, integrating interpretability methods into workflow-specific interfaces makes it possible to study their performance within tailored tasks. For example, Wu et al. (2021) designed an interface to aid auditors in generating counterfactual examples that flip model classifications, helping to characterize sensitivity to out-of-distribution shifts. Influential examples that participants discovered could be incorporated into training to improve generalization. Different choices in the frontend interface and backend interpretability engine led to differential success in discovering surprising counterfactuals (those with small perturbations in input that led to large change in class predictions). In this way, interface-oriented user studies can more precisely gauge “human-AI collaboration.” Indeed, these experiments have uncovered counterintuitive phenomena, like instances where task performance worsens with AI explanations — human “collaborators” can become less critical when shown an explanation (Bansal et al., 2021). Finally, a key challenge in user studies is selecting an appropriate participant pool with representative proxy tasks (Buçinca et al., 2020). For example, an interpretability method in healthcare may be more effective for patients than for doctors. Narrow tasks can more directly guide deployment while broader ones are easier to run at scale.
@@
There is often overlap between benchmark datasets and user study evaluation. For example, Casper et al. (2023) created a dataset to understand the effectiveness of explanations in aiding model debugging. The dataset introduced distractor patches to a small fraction of images, leading to failures on the test set. For example, a small strawberry was pasted into 1 in 3000 images in both training and test sets, and in each “poisoned” training image, the class was switched to “goose,” leading to failures at test time. Participants were presented with misclassified test images, along with various explanations. Their accuracy and speed in identifying the poisoning mechanism served to quantify the utility of these explanations for debugging.
@@
Evidently, while there is no one way to evaluate interpretability methods, there is a natural hierarchy of tests to which proposals can be subjected. General-purpose proposals should first pass ablation and synthetic data benchmarks. Success in these criteria justifies investing resources in user studies with more realistic audience and task designs. Future interpretability methods may be evaluated similarly to new therapies in medicine: in silico associations (ablation and synthetic data studies) can guide in vitro experiments (user studies with generic audiences) which, if successful, lead to follow-up clinical trials (user studies with representative audiences).
@@ 4 Discussion
This review has surveyed the significant advances in interpretability over recent decades. In closing, we highlight key ongoing shifts in the AI landscape and their implications for interpretability. One important trend is the rise of multimodal models, which are trained on mixed data modalities (Ngiam et al., 2011; Chen et al., 2021). This has facilitated novel applications, like automatic open vocabulary segmentation (Zou et al., 2023), where, rather than annotating each image pixel with one of a preset collection of class labels, the system can generate segmentation masks for arbitrary text strings (e.g., “The pedestrian that is crossing the street.”). While single modality methods remain applicable to individual model components (e.g,. analyzing the word embeddings in a vision language model), there is a pressing need for interpretability techniques that offer an integrated view of how these models merge information across sources. This could help developers train models that effectively balance evidence from complementary views, enhancing overall system robustness. Statistical methods for data integration (Ding et al., 2022) could offer a principled approach for analyzing models with these more complex, interacting components.
@@
Third, as new legal and social norms around algorithmic decision-making are established, we will likely encounter scenarios where even comprehensive explanation are unsatisfactory, and where high-performing glass boxes become critical. In certain applications, reliability and interpretability will outweigh performance alone. For example, in healthcare and legal contexts, simple decision-making criteria are ideal, and in science, generalizable insight holds greater value than out-of-sample performance alone. Though recent progress in deep learning has sparked the community’s imagination about potentially more human-like or human-surpassing systems, we should perhaps consider futures where systems only marginally outperform current ones but are significantly safer and transparent.
Finally, we ask: how can we make interpretability a more interactive process? Interactivity is critical for ensuring that our systems augment intelligence, empowering users to become creative and critical problem solvers. Neglecting this problem may lead to an overreliance on oracles, even among experts. For example, doctors should feel comfortable rejecting AI recommendations, even if there is a chance their decision turns out to be incorrect. A prerequisite for such interactivity is shared language (Heer, 2019; Crabbe et al., 2021; Kim, 2022) – it is difficult to interact without communication. CBMs are already a step in this direction, where concepts serve as a representation linking both human and machine reasoning. The goal of a shared language isn’t as fantastical as it might seem; programming, after all, is an exercise in communicating with machines, with the programming language as the shared representation.
@@
Acknowledgement
This work was supported by award R01GM152744 from the National Institute of General Medical Sciences of the National Institutes of Health.
We also thank the organizers and participants of the BIRS Workshop, “Statistical Aspects of Trustworthy Machine Learning” for creative and critical discussion.

### v2

@@ 3 Evaluation
Human-Computer Interaction: The success of an interpretability method depends on how it interfaces with users (Baniecki et al., 2021; Baniecki and Biecek, 2019). Simplified experiments may not reflect real-world complexity, while real-world observations may be confounded by factors that don’t generalize across applications.
@@
Interpretability evaluation methods fall into two general categories: tailored benchmark datasets and user studies with controlled tasks. Within each class, there are variations that account for the challenges above in different ways. For dataset benchmarks, one approach is ablation (Hooker et al., 2019; Wang et al., 2022). Specifically, for methods that output feature importance attributions, this involves retraining models with versions of the data where purported important regions are masked out. Deterioration in model performance suggests that the attributions accurately reflect regions that the original model relied upon. Note that retraining is necessary. The ablated images are not drawn from the same distribution as the training data, so deterioration in performance of the original model could be due to either removal of important features (the criteria of interest) or domain shift (a confounding factor). Another data-driven benchmarking approach is to design synthetic datasets. These datasets are generated using known mechanisms, allowing for later evaluation of explanations (Liu et al., 2021; Zhou et al., 2022). For example, the funnybirds benchmark creates imaginary cartoon birds where “species” classes are derived from attributes, like wing, beak, and tail type, that are controlled during data generation (Hesse et al., 2023). Given a specific instance, an ideal explanation should focus on the parts of the bird that distinguish it from related classes. To make these benchmarks more realistic, variation in viewpoint and illumination can be introduced.
@@
In user studies, one approach is to have participants make pairwise preference judgments between explanations (Jeyakumar et al., 2020). However, subjective preference may not align with objective task performance. To this end, more specific experimental tasks can be designed, evaluating user decisions with and without support from interpretability outputs (Ma et al., 2024; Colin et al., 2022). For example, to assess simulatability, participants could be asked to guess a model’s output on an example before and after viewing an explanation. Alternatively, they could be asked to guess how predictions might change under perturbations to the input – effective explanations should help in making accurate guesses. Finally, integrating interpretability methods into workflow-specific interfaces makes it possible to study their performance within tailored tasks. For example, Wu et al. (2021) designed an interface to aid auditors in generating counterfactual examples that flip model classifications. This is an instance of a counterfactual explanation (Mothilal et al., 2020; Guidotti, 2022) that helps to characterize sensitivity to out-of-distribution shifts. Influential examples that participants discovered could be incorporated into training to improve generalization. Different choices in the frontend interface and backend interpretability engine led to differential success in discovering surprising counterfactuals (those with small perturbations in input that led to large change in class predictions). In this way, interface-oriented user studies can more precisely gauge “human-AI collaboration.” Indeed, these experiments have uncovered counterintuitive phenomena, like instances where task performance worsens with AI explanations — human “collaborators” can become less critical when shown an explanation (Bansal et al., 2021). Finally, a key challenge in user studies is selecting an appropriate participant pool with representative tasks (Buçinca et al., 2020). For example, an interpretability method in healthcare may be more effective for patients than for doctors. Narrow tasks can more directly guide deployment while broader ones are easier to run at scale.
@@
There is often overlap between benchmark datasets and user study evaluation. For example, Casper et al. (2023) created a dataset to understand the effectiveness of explanations in aiding model debugging. The dataset introduced distractor patches to a small fraction of images, leading to failures on the test set. For example, a small strawberry was pasted into 1 in 3000 images in both training and test sets, and in each “poisoned” training (but not test) image, the class was switched to “goose,” leading to failures at test time. Participants were presented with misclassified test images, along with various explanations. Their accuracy and speed in identifying the poisoning mechanism served to quantify the utility of these explanations for debugging.
@@
Evidently, while there is no one way to evaluate interpretability methods, there is a natural hierarchy of tests to which proposals can be subjected. General-purpose proposals should first pass ablation and synthetic data benchmarks. Success in these criteria justifies investing resources in user studies with more realistic audience and task designs (Rong et al., 2024). Future interpretability methods may be evaluated similarly to new therapies in medicine: in silico associations (ablation and synthetic data studies) can guide in vitro experiments (user studies with generic audiences) which, if successful, lead to follow-up clinical trials (user studies with representative audiences).
@@ 4 Discussion
This review has surveyed the significant advances in interpretability over recent decades. In closing, we highlight key ongoing shifts in the AI landscape and their implications for interpretability. One important trend is the rise of multimodal models, which are trained on mixed data modalities (Ngiam et al., 2011; Chen et al., 2021). This has facilitated novel applications, like automatic open vocabulary segmentation (Zou et al., 2023), where, rather than annotating each image pixel with one of a preset collection of class labels, the system can generate segmentation masks for arbitrary text strings (e.g., “The pedestrian that is crossing the street.”). While single modality methods remain applicable to individual model components (e.g., analyzing the word embeddings in a vision language model), there is a pressing need for interpretability techniques that offer an integrated view of how these models merge information across sources. This could help developers train models that effectively balance evidence from complementary views, enhancing overall system robustness. Statistical methods for data integration (Ding et al., 2022) could offer a principled approach for analyzing models with these more complex, interacting components.
@@
Third, as new legal and social norms around model-based decision-making are established, we will likely encounter scenarios where even comprehensive explanation are unsatisfactory, and where high-performing glass boxes become critical (Nannini et al., 2023). In certain applications, reliability and interpretability will outweigh performance alone. For example, in healthcare and legal contexts, simple decision-making criteria are ideal, and in science, generalizable insight holds greater value than out-of-sample performance alone. Though recent progress in deep learning has sparked the community’s imagination about potentially more human-like or human-surpassing systems, we should perhaps consider futures where systems only marginally outperform current ones but are significantly safer and transparent.
Finally, we ask: how can we make interpretability a more interactive process? Interactivity is critical for ensuring that our systems augment intelligence, empowering users to become creative and critical problem solvers (Baniecki et al., 2023; Sokol and Flach, 2020). Neglecting this problem may lead to an overreliance on oracles, even among experts. For example, doctors should feel comfortable rejecting AI recommendations, even if there is a chance their decision turns out to be incorrect. A prerequisite for such interactivity is shared language (Heer, 2019; Crabbe et al., 2021; Kim, 2022) – it is difficult to interact without communication. CBMs are already a step in this direction, where concepts serve as a representation linking both human and machine reasoning. The goal of a shared language isn’t as fantastical as it might seem; programming, after all, is an exercise in communicating with machines, with the programming language as the shared representation.
@@
Supplementary Material
Code to reproduce our simulation experiment can be found at https://go.wisc.edu/3k1ewe. The repository’s README file describes how to generate the dataset, fit each model, and create the visualizations. The data, compiled notebooks, intermediate outputs, and log files can be accessed at https://go.wisc.edu/v623lq.
Funding
This work was supported in part by grant number R01GM152744 from the National Institute of General Medical Sciences.